solve x²(y-z)p+y²(z-x)q=z²(x-y)=0


1
Expert's answer
2021-03-16T04:56:31-0400

Given the equation:


"x\u00b2(y-z)p+y\u00b2(z-x)q-z\u00b2(x-y)=0"

We can rewrite the equation as:


"x\u00b2(y-z)p+y\u00b2(z-x)q=z\u00b2(x-y)"

The above equation is of the form:


"Pp +Qq = R"

Where:


"P = x²(y-z)\\
Q=y²(z-x)\\
R= z²(x-y)"

The auxilliary equation is of the form:


"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

Substituting for P, Q and R:


"\frac{dx}{x²(y-z)}=\frac{dy}{y²(z-x)}=\frac{dz}{z²(x-y)} \qquad \cdots eqn (i)"



Solvijng the above:


"\frac{\frac{dx}{x²}}{y-z}=\frac{\frac{dy}{y²}}{z-x}=\frac{\frac{dz}{z²}}{x-y}"

"\frac{\frac{dx}{x²}+\frac{dy}{y²}+\frac{dz}{z²}}{y-z+ z-x+x-y}=k \qquad \text{(say)}\\
\frac{\frac{dx}{x²}+\frac{dy}{y²}+\frac{dz}{z²}}{0}=k\\
\implies \frac{dx}{x²}+\frac{dy}{y²}+\frac{dz}{z²}=0"

Integratiing through:


"\int \frac{dx}{x²}+\int \frac{dy}{y²}+\int \frac{dz}{z²}=\int 0\\
-\frac{1}{x}-\frac{1}{y}-\frac{1}{z}= c\\
\implies c_1 = -\Big(\frac{1}{x}+\frac{1}{y}+\frac{1}{z} \Big)"

Also,


"\frac{\frac{dx}{x}}{x(y-z)}=\frac{\frac{dy}{y}}{y(z-x)}=\frac{\frac{dz}{z}}{z(x-y)}\\
\frac{-\frac{dx}{x}}{x(z-y)}=\frac{-\frac{dy}{y}}{y(x-z)}=\frac{-\frac{dz}{z}}{z(y-x)}\\
\frac{-\frac{dx}{x} + (-\frac{dy}{y})+ (-\frac{dz}{z})}{x(z-y)+y(x-z)+z(y-x)} = k \qquad \text{(say)}\\
\frac{-\frac{dx}{x} -\frac{dy}{y}-\frac{dz}{z}}{xz-xy+xy-yz+yz-xy} = k\\
\frac{-\frac{dx}{x} -\frac{dy}{y}-\frac{dz}{z}}{0} = k\\
\implies -\frac{dx}{x} -\frac{dy}{y}-\frac{dz}{z} = 0"

Integrating through:


"\int-\frac{dx}{x}\int -\frac{dy}{y}\int-\frac{dz}{z}=\int 0\\
-\int\frac{1}{x}dx-\int \frac{1}{y}dy-\int\frac{1}{z}dz=\int 0\\
- \ln x - \ln y - \ln z = \ln c\\
\ln x^{-1} + \ln y^{-1} + \ln z^{-1} = \ln c\\
\ln \Big(\frac{1}{x}\Big) + \ln \Big(\frac{1}{y}\Big) + \ln \Big(\frac{1}{z}\Big) = \ln c\\
\ln \Big(\frac{1}{xyz} \Big) =\ln c\\
\frac{1}{xyz} = c\\
\implies c_2 = \frac{1}{xyz}"

Therefore, the general solution is:


"f(u,v) = \Bigg(-\Big(\frac{1}{x}+\frac{1}{y}+\frac{1}{z} \Big), \frac{1}{xyz} \Bigg) = 0"